OpenStudy (anonymous):

If a @ b = 2(a + |b|), what is the value of 2 @ (-2 @ -1)? A. -8 B. 4 C. 0 D. 4 E. 8 wat should i do after this point? -2 @ -1 = 2(-2 + |-1|) = 2(-2 + 1) = 2(-1) = -2

OpenStudy (hba):

2@-2=2(2+2)=8

OpenStudy (anonymous):

so like where did u get the 2(2+2) part again?

OpenStudy (lgbasallote):

i think you need to solve foe -2 @ -1 first

OpenStudy (lgbasallote):

-2 @ -1 = 2(-2 + |-1|)

OpenStudy (lgbasallote):

then substitute that 2 @ 2(-2 + |-1|)

OpenStudy (lgbasallote):

now use a @ b again

OpenStudy (lgbasallote):

2(2 + |2(-2 + |-1|)|)

OpenStudy (hba):

a@b=2(a+|b|)=2@-2=2(2+|-2|)=2(2+2)

OpenStudy (lgbasallote):

-2 + |-1| = -2 + 1 = -1 2(2 + |2(-1)|) 2(2 + |-2|) 2(2 + 2) 2(4)

OpenStudy (lgbasallote):

i hope i made sense lol

OpenStudy (anonymous):

Lol im still kind of loading i take a lil bit of time on these qsts...

OpenStudy (anonymous):

ok i think i just got it and i found the answer to be 8

OpenStudy (lgbasallote):

yup